OpenStudy (anonymous):

How many distinct natural numbers are in the solution set of {x:|2x+6|≥4} and (intersection) {x:|3x+20|≤29} ? Please explain.

OpenStudy (anonymous):

I'm not sure about this notation. Are those absolute value signs?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Start with |2x + 6| >= 4 is 2 sets, the first of which, for 2x + 6 greater than 0 -> 2x +6 >= 4 and the second, for 2x + 6 less than 0, 2x + 6 <= -4.

OpenStudy (anonymous):

yeah i know that the first equation splits in x≥-1 or x≤-5 and the second equation splits into x≤3 and x≥-49/3 i just don't know how to find how many natural numbers there are in the solution

OpenStudy (anonymous):

For |3x + 20| <= 29, you have -29 <= 3x + 20 <= 29

OpenStudy (anonymous):

I'd say it's infinity

OpenStudy (anonymous):

You can list each of the "halves" with ellipses notation (...) after solving and then work the intersection.

OpenStudy (anonymous):

just kidding

OpenStudy (anonymous):

okay so do you know the solution set of the two equations?

OpenStudy (anonymous):

So, working with the second set, or -29 <= 3x + 20 <= 29 we get -49 <= 3x <= 9 or (-49)/3 <= x <= 3. So this will give the explicit values for x for the SECOND set, once you take into account only natural numbers. Remember your definition of natural numbers.

OpenStudy (anonymous):

yes you do. now the natural numbers are from 1 on to infinity (0 is disputed) so you can discard all the negative numbers in the solution sets. the intersection of the positive solution sets is you answer

OpenStudy (anonymous):

can you help me find the intersection of the positive solution sets?

OpenStudy (anonymous):

Remeber that natural numbers are 1,2,3,... That is the "counting" numbers. Starting with "1" and no negatives. "0" is not disputed in the set of natural numbers. If you include "0", you get the "whole" numbers.

OpenStudy (anonymous):

wait is the answer 3?

OpenStudy (anonymous):

one of the solutions sets is -1to infinity and the other one is x less than equal to 3. yes

OpenStudy (anonymous):

The solution set to to the second inequality is [1,2,3]. Now, we have to look again at the first inequality.

OpenStudy (anonymous):

it's 3

OpenStudy (anonymous):

We do not want negative numbers, so for the first inequality, we just have to look at 2x +6 >= 4. That set will be all natural numbers. So the intersection set is {1,2,3}. Three numbers. Done.